OpenStudy (anonymous):

solve (n-2/n+6)=(1/5)

OpenStudy (anonymous):

anyone?

OpenStudy (abb0t):

\[\frac{ n-2 }{ n+6 }= \frac{ 1 }{ 5 }\] 1. multiply the fraction 1/5 by (n+6). 2. Distribute 1/5 into the parenthesis You should get: \[n-2 = \frac{ 1 }{ 5 }n+\frac{ 6 }{ 5 }\] 3. add 2 to both sides (this cancels out the 2 on the left. Now you have to add (2 + 6/5) 4. Subtract 1/5n from n. At this point you should have #n = # 5. now solve for n by dividing and isolating n by itself. You should be dividing the two numbers to get n = #/# = # where # = some number that you got.

OpenStudy (anonymous):

n= 6 n= 5 n= 4 these are my options

OpenStudy (abb0t):

Try solving for it as I explained. I already gave you the first step. Also, it might help to know this property of fractions: \[\frac{ 1 }{ b }n= \frac{ 1 }{ b } \] means you are dividing to solve for n. Therefore:\[\frac{ 1 }{ b }n = \frac{ \frac{ 1 }{ b } }{ \frac{ 1 }{ b } }= \frac{ 1 }{ b } \times \frac{ b }{ 1 } = 1\]